OpenStudy (anonymous):

find the limit:

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \sin 3x/2x\] Also i know that sinx/x=1 is that relevant

OpenStudy (anonymous):

yes, but do you know why lim sinx/x=1? That will get you there.

OpenStudy (anonymous):

i don't i just remember thats a special property

OpenStudy (anonymous):

a nudge in the right direction?

OpenStudy (kropot72):

You can find the limit by as follows: 1) Express sin3x as an expansion in series up to say the fifth term. 2) Divide each term of the series expansion by 2x. 3) You can easily find the limit at this point

OpenStudy (anonymous):

i'm not sure i understand about series. what do you mean?

OpenStudy (kropot72):

The series expansion of f(x) can be found from Maclaurin's theorem:\[\sin x=x-\frac{x ^{3}}{3!}+\frac{x ^{5}}{5!}-\frac{x ^{7}}{7!}+....\]

OpenStudy (anonymous):

i've never learned any of this

OpenStudy (anonymous):

is there another way to do it?

OpenStudy (anonymous):

nevermind i'l just skip these questions